# Leveraging Large Language Models for Collective Decision-Making

### v2

@@ 3.2. Coordination Process
The discussion aims to find an option for the group decision that maximally meets the members’ preferences through many rounds of interactions between the members and the system. Toward this goal, we define decision candidate for each round $t$ as the option from the set of option $\mathcal{X}^{t}$ that satisfies most of the members; we denote it as $\hat{x}^{t}\in\mathcal{X}^{t}$. For the coordination process to work, we assume that there is always at least one option that satisfies at least two members.
We describe how the decision candidates from previous rounds are incorporated into the latest round in the algorithm as below. At $t=0$, we initialize $\mathcal{X}^{0}=\emptyset$ and $\hat{x}^{0}=\emptyset$. Figure 1(b) illustrates the coordination algorithm for round $t\in\{1,2...T\}$, as detailed in the following steps:
@@
    The system presents the set of options based on round $t-1$, i.e., $\mathcal{X}^{t-1}$, to each member, Observing $\mathcal{X}^{t-1}$, each member $i$ has their conversation $m_{i}^{t}$ with the system.
@@
    The intent extraction module receives the conversation $\bm{m}^{t}$ and outputs the extracted preferences $\bm{\theta}^{t}$ to the coordinator module.
@@
    Based on the extracted preferences $\bm{\theta}^{t}$ (and the topic and knowledge graph $\mathcal{G}$), the coordinator produces a set of $K$ options $\mathcal{X}^{t}$ as follows:
    • For $t=1$, as there is no candidate decision from the previous round, i.e., $\hat{x}^{0}=\emptyset$, the coordinator module suggests $K$ options, each of which should satisfy at least 2 members.
    • For $t>1$, there is a decision candidate from round $t-1$ as constructed in step (4) (i.e., $\hat{x}^{t-1}\neq\emptyset$), the coordinator module presents $\hat{x}^{t-1}$ and proposes another $K-1$ options that should satisfy at least as many members as the candidate decision $\hat{x}^{t-1}$ does.
    Each option $x_{j}^{t}\in\mathcal{X}^{t}$ contains three fields for informational purposes: (i) suggestion for the group decision, (ii) the subset of the members this option satisfies (i.e., any $i$ such that $\pi_{j,i}^{t}>0$), and (iii) a list of reasons for this suggestion.
@@
    Namely, the system first identifies the set of options that satisfy the greatest number of members and then selects an option with the largest satisfaction score. The rationale behind the definition of the decision candidate is the following: In constructing $\mathcal{X}_{\star}^{t}$, we identify the option that satisfies as many members as possible, i.e., we minimize the number of the members whose preferences are not fulfilled at all by the option. Then, in case there are ties between the potential decision candidates in $\mathcal{X}_{\star}^{t}$, we keep the option that fulfills the largest number of the preferences of these members, i.e., the largest satisfaction score. [^2] If multiple options still have the same maximum satisfaction score, the ties are broken arbitrarily, and one decision candidate is selected. Note that the decision candidate is replaced with a new option if the new option yields a higher satisfaction ratio or satisfaction score.
For the remainder of the paper, we use $\hat{r}^{t},\hat{s}^{t},\hat{g}^{t}$ to denote the satisfaction ratio, satisfaction score, and equity score of decision candidate $\hat{x}^{t}$, respectively.
@@ 4. A Case Study: Group Meeting Scheduling
We show an example scenario of meeting scheduling in Figure 3, where an organizer [^3] In the context of meeting scheduling, we use the term “organizer” to refer to the group leader and use the “meeting scenarios” for the group discussions. wants to schedule a meeting for $n=3$ members. The meeting organizer sends out a meeting invite via the system (Figure 3(a)) that describes the meeting’s purpose. Next, each member converses with the intent extraction modules during which their preferences are solicited (see Figures 3(b), 3(c) and 3(d)). The collected preferences are then sent to the coordination module for processing.
After processing the preferences, the coordination module proposes $K=2$ options as shown in Figure 3(e) and presents them to the members. The first option is a meeting in the morning that meets all of the revealed preferences of Member 1 and Member 2. The second option is an afternoon meeting, which satisfies all of Member 2’s preferences and more than half of Member 3’s preferences.
@@
To implement our system for meeting scheduling leveraging LLMs, we designed a set of four prompts for the modules using gpt-3.5-turbo. The details of the prompts can be found in Appendix B. In particular, the evaluation module was designed to assign a value to any given member $i$ and option $x_{j}^{t}$ with Likert-scale-like values:
(5) $\pi_{j,i}^{t}=\begin{cases}0,&\text{if option $j$ satisfies none of the preferences in $\theta_{i}^{t}$};\\ 1,&\text{if option $j$ satisfies at most half of the preferences in $\theta_{i}^{t}$};\\ 2,&\text{if option $j$ satisfies at least half but not all of the preferences in $\theta_{i}^{t}$};\\ 3,&\text{if option $j$ satisfies all of the preferences in $\theta_{i}^{t}$}.\end{cases}$
The satisfaction score $s_{j}^{t}$, which is derived based on $\{\pi_{j,i}^{t}\}_{i\in[n]}$, in turn lies between 0 and 3. We use this discretized scale instead of fractions (i.e., range of 0 to 1) since we did a test and empirically observed that LLMs were less capable of dealing with fractions and errors often occurred.
We assess the performance of leveraging LLMs to coordinate meeting scheduling. We study various aspects during coordination dynamics, including the efficiency of eliciting preferences, the effectiveness of identifying mutually agreed-upon options, the satisfaction of members’ preferences, and the quality of reasoning and preference aggregation by LLMs. Specifically, we aim to answer the following research questions (RQs):
@@
A natural way to evaluate the system and answer (RQ1)-(RQ4) is by recruiting human participants and conducting a study in which the human participants use the system for meeting scheduling. However, this approach presents certain challenges in terms of implementation and scalability. Precisely, to ensure the significance of the study’s results, it is necessary to conduct numerous meeting scheduling scenarios. Since each scenario involves several rounds of conversations between human participants and the system, such a study requires many human participants and time to complete. For instance, the user study of Calendar.help (Cranshaw et al., 2017) lasted for 18 months. We believe that enabling a user study to be completed quickly would be beneficial to iterate over and expedite system designs.
Thus, to address these challenges, we alternatively conducted the evaluation study under synthetic settings of meeting scheduling and used LLMs to simulate members’ conversation behavior. This approach enables more controlled and scalable experimentation while preserving a realistic level of human-like interaction.
@@ 5.1. Synthetic Data
We created a synthetic medical start-up company with 34 employees. Each employee was assigned six fields of information: (1) employee name, (2) a role (e.g., CEO, Software Engineer, Sales Manager, etc.), (3) their manager, (4) a seniority level ranging from 1 to 5, (5) responsibilities (e.g., “Assisting the sales team in various tasks and projects”), and (6) schedule preferences (e.g., prefer to have meetings in the afternoon, to have time to drop the kids at school in the morning, etc.). Table 4 shows some example employee data in the aforementioned structure.
We synthesized data of meetings that includes: (1) an organizer, (2) members of the meetings, (3) the meeting subject, (4) date, and (5) duration. Example meeting data can be found in Table 5.
@@
Finally, the knowledge graph $\mathcal{G}$ for this synthetic company encompasses all the metadata about employees that is accessible to the system. Precisely, the knowledge graph includes the social graph and the profile data of all employees except for their schedule preferences (which are their private information and should be elicited by the system).
In the sequel, we detail our experiment setup, evaluation metrics, and results.
@@ 5.2. Experiment Setup
We evaluated the performance of our system on the following axes: (i) the effect of the parameters of the system ($n,K$, incorporation of the knowledge graph) on the performance, (ii) system performance compared to two baselines, (iii) robustness of our system subject to paraphrased preferences.
@@
  • Schedule preferences only. First, the system was required to propose meeting options based solely on the employees’ schedule preferences without considering the knowledge graph $\mathcal{G}$. Namely, the system focused on finding a meeting time that balances individual schedule preferences elicited from meeting members. For this setting, we examined the effect of the meeting size $n$ and the number of proposed options $K$ as follows:
@@
  • Schedule preferences and Knowledge Graph. Second, the coordination module considered not only schedule preferences but also the information embedded in the knowledge graph $\mathcal{G}$ (i.e., the social subgraph induced on the meeting members, and their roles and responsibilities); the system also accounted for attendance constraints (if any) posed by the meeting organizer.
    To examine the effect of incorporating the knowledge graph in the system’s suggestions, we tested the system on 20 meeting scenarios with $n=5$ and $K=2$. First, we focused on $n=5$ members since in smaller meetings, the structure of the social subgraph was not complex enough to effectively showcase the impacts of the social graph on the systems’s reasoning. Furthermore, based on the findings of the baseline study (cf. Section 6), the system’s performance metrics deteriorated with larger values of $K$, and, therefore, we opted for $K=2$.
In all of simulated meeting scenarios, the coordination processes were carried out over a total of $T=4$ rounds. This is consistent with the finding of Calendar.help (Cranshaw et al., 2017) which shows that most users had no more than five rounds of interactions with the system.
@@
In total, we ran 120 meeting scenarios. We note that throughout the two studies, we chose $n>K$, i.e., the meeting size was always greater than the number of proposed options. This is because when the options presented at each round are more than the meeting members, it is difficult to amass a majority on one of the options and, therefore, to reach a consensus. Moreover, the coordination process is designed with the goal of collecting the members’ feedback and gradually increasing the fraction of the members that the decision candidate can satisfy (i.e., satisfaction ratio $\hat{r}^{t}$). Therefore, the decision candidate is refined via several rounds of proposing new options, and proposing many options at each round is not needed.
Comparison with baselines. Our system includes two interactive elements: (i) the conversation between a user and the intent extraction module, (ii) the multiple rounds of discussion. To show the gains from the two elements, we compared the system with the following two baselines: [^5] We remark that since our proposed system focuses on open-ended textual preferences, a direct comparison with the methods that process on time slots (e.g., Doodle, when2meet, MS Outlook, Calendar.help etc.) is not viable.
  • Single-Round Non-Conversational Baseline. In this baseline, both elements were absent. We did not use the interactive/conversational intent extraction module and instead provided the raw preferences $\bm{\theta}^{1}$ directly to the coordination module. We ran the protocol for only one round (i.e., $T=1$).
  • Single-Round Conversational Baseline. In this baseline, we used the interactive intent extraction module for the conversation with users, but ran the system for only one round. This is to measure the extent to which multiple rounds of refining the decision candidate improves the system performance.
@@ 6. Simulation Study Results
In this section, we present the results of our study in Section 5 and Section 7. Figures 6, 4 and 5 summarize the evaluation metrics in the experiment of Section 5, which are plotted in the following manner: Given a meeting scenario, for each metric we calculated its percentage change at round $t$ with respect to its value at round $t=1$. Then we plotted the average of the percentage change in the metrics for each round, together with their 95% confidence intervals in our system [^6] We use the Student-t distribution with 19 degrees of freedom and a two-sided confidence interval.. The plots showing the metrics values for all scenarios are relegated to Appendix C.
Figure 4. We plot the average of the percentage change in the metrics with the number of rounds for the settings $T=4,K=5$ and $n\in\{3,4,5\}$. For each setting, 20 different meeting scenarios were used. Given the metrics of a meeting scenario, the percentage change for each round is computed with respect to $t=1$. The plots show the average of the percentage changes over the 20 scenarios. The legends in the top-left corner of each sub-figure correspond to the average initial value and the average final value of each metric, respectively.
@@ 6.1. Effect of Meeting Size $n$
Figure 4 shows the results when meeting size $n$ varies. We observed three trends across the three settings examined: (i) the average number of iterations is decreasing, (ii) the satisfaction ratio is increasing, (iii) the satisfaction score is increasing.
In particular, the system performed significantly better in the setting with $n=5$ members than the ones with $n=3$ and $n=4$ in these three dimensions. For $n=5$ members, the average number of interactions was largely reduced by $\approx 50\%$ [^7] We use $\approx$” to for “approximately”., compared to $\approx 25\%$ ($n=3$) and $\approx 45\%$ ($n=4$). Secondly, the satisfaction score of the decision candidate for $n=5$ was eventually increased by $\approx 120\%$, while it was increased by $\approx 45\%$ ($n=3$) and $\approx 40\%$ ($n=4$) respectively. Similarly, the satisfaction ratio of the decision candidate was increased by $\approx 110\%$ for $n=5$, while both $n=3$ and $n=4$ experienced an increase of $\approx 40\%$. Finally, the equity score of the decision candidate decreased for $n=5$, but it exhibited subtle fluctuations for $n=3$ and $n=4$.
@@ 6.2. Effect of Number of Options $K$
Figure 5 shows the results of the experiment for the settings when the number of options $K$ varies. We found that when the number of options was increased from $K=2$ to $K=4$, there was a decrease in the final percentage change of the satisfaction ratio. Specifically, for $K=2$ options, the system achieved an $\approx 100\%$ increase in the satisfaction ratio at the final round, whereas adopting $K=3$ and $K=4$ options yielded an increase of $\approx 50\%$. We also found that the increase in satisfaction score was $\approx 100\%$ for $K=2$, compared to $\approx 50\%$ ($K=3$) and $\approx 100\%$ ($K=4$). The equity score experienced a smaller percentage decrease as $K$ was increased. These observations indicated that the system’s performance was negatively impacted as the number of options $K$ was increased.
Figure 6. The average of the percentage change in each metric with the number of rounds for $T=4,n=5,K=2$ and whether the knowledge graph $\mathcal{G}$ is considered or not. For each setting, 20 different meeting scenarios were tested.
6.3. Effect of Incorporating the Knowledge Graph
Figure 6 compares the metrics between including and excluding the knowledge graph for $T=4$, $n=5$ and $K=2$. When the knowledge graph and attendance constraints were incorporated, the average number of interactions was still decreasing, and both of the satisfaction ratio and the score of the decision candidate were increasing. However, compared to using the schedule preferences only, the evaluation metrics deteriorated, and the system’s performance was worse in each dimension.

### v3

@@ 3.2. Coordination Process
The system assists the group in identifying and revising their decision to meet the members’ preferences as much as possible through many rounds of private conversations between the members and the system. Toward this goal, we define decision candidate for each round $t$ as an option from the set of option $\mathcal{X}^{t}$ that satisfies most of the group members; we denote it as $\hat{x}^{t}\in\mathcal{X}^{t}$.
Below, we describe the system protocol of how the group decision is found and improved through the $T$ rounds, as illustrated in Figure 1(b). At $t=0$, we initialize $\mathcal{X}^{0}=\emptyset$ and $\hat{x}^{0}=\emptyset$. At round $t\in\{1,2...T\}$,
@@
    The system presents to each member the set of options suggested based on round $t-1$, i.e., $\mathcal{X}^{t-1}$.
@@
    Observing $\mathcal{X}^{t-1}$, each member $i$ has their private conversation $m_{i}^{t}$ with the intent extraction module, and the intent extraction module outputs the extracted preferences $\theta_{i}^{t}$ to the coordination module.
@@
    According to the collection of members’ preferences $\bm{\theta}^{t}$, the group’s intent, and relevant information from database $\mathcal{D}$, the coordinator generates a set of $K$ options $\mathcal{X}^{t}$ as follows:
    • For $t=1$, no candidate decision has been proposed for the group (i.e., $\hat{x}^{0}=\emptyset$). Therefore, the coordinator module is instructed to suggest $K$ options, and each of the options should satisfy at least 2 members.
    • For any $t>1$, there is a decision candidate from round $t-1$ as constructed in step (4) (i.e., $\hat{x}^{t-1}\neq\emptyset$), the coordinator module presents $\hat{x}^{t-1}$ and proposes another $K-1$ options that should satisfy at least as many members as the candidate decision $\hat{x}^{t-1}$ does.
@@
    Namely, the system first identifies the set of options that satisfy the greatest number of members and then selects an option with the largest satisfaction score. We design this selection rule for the following reasons: By constructing $\mathcal{X}_{\star}^{t}$, we minimize the number of members whose preferences are not fulfilled at all by the option. Then, in case there are ties between the potential decision candidates in $\mathcal{X}_{\star}^{t}$, we choose the option that maximizes satisfaction score, i.e., fulfills a large number of the preferences of these members. [^4] If multiple options still have the same maximum satisfaction score, the ties are broken arbitrarily, and one decision candidate is selected. Note that the decision candidate is replaced with a new option if the new option yields a higher satisfaction ratio or satisfaction score.
For the remainder of the paper, we use $\hat{r}^{t},\hat{s}^{t},\hat{g}^{t}$ to particularly denote the satisfaction ratio, satisfaction score, and equity score of decision candidate $\hat{x}^{t}$, respectively.
@@ 4. A Case Study: Group Meeting Scheduling
We show an example scenario of meeting scheduling in Figure 3, where an organizer wants to schedule a meeting for $n=3$ members. [^5] In the context of meeting scheduling, we use the term “organizer” to refer to a member who leads the group decision process and use the “meeting scenarios” for the group decisions. The meeting organizer sends out a meeting invite via the system (Figure 3(a)) that describes the meeting’s purpose. Next, each member converses with the intent extraction modules during which their preferences are solicited (see Figures 3(b), 3(c) and 3(d)). The collected preferences are then sent to the coordination module for processing.
[…]
@@
To implement our system for meeting scheduling leveraging LLMs, we designed a set of four prompts for the modules using gpt-3.5-turbo [^6] The experiments have been run between June and August 2023.. The details of the prompts can be found in Appendix B. In particular, the evaluation module was designed to assign a value to any given member $i$ and option $x_{j}^{t}$ with Likert-scale-like values:
(5) $\pi_{j,i}^{t}=\begin{cases}0,&\text{if option $x_{j}^{t}$ fulfills none of the preferences in $\theta_{i}^{t}$};\\ 1,&\text{if option $x_{j}^{t}$ fulfills at most 50\% of the preferences in $\theta_{i}^{t}$};\\ 2,&\text{if option $x_{j}^{t}$ fulfills at least 50\% but not all of the preferences in $\theta_{i}^{t}$};\\ 3,&\text{if option $x_{j}^{t}$ fulfills all of the preferences in $\theta_{i}^{t}$}.\end{cases}$
The satisfaction score $s_{j}^{t}$, which is derived based on $\{\pi_{j,i}^{t}\}_{i\in[n]}$, in turn lies between 0 and 3. We use this discretized scale instead of fractions (i.e., range of 0 to 1) since we did a test and empirically observed that LLMs were less capable of dealing with fractions, and errors often occurred.
We assess the performance of leveraging LLMs to coordinate meeting scheduling. We study various aspects of coordination dynamics, including the efficiency of eliciting preferences, the effectiveness of identifying mutually agreed-upon options, the satisfaction of members’ preferences, and the quality of LLMs’ reasoning and preference aggregation.
Specifically, we aim to answer the following research questions (RQs):
@@
A natural way to evaluate the system and answer (RQ1)-(RQ4) is by recruiting human participants and conducting a study in which the human participants use the system for meeting scheduling. However, this approach presents certain challenges in terms of implementation and scalability. To ensure the significance of the study’s results, it is necessary to conduct numerous meeting scheduling scenarios. Since each scenario involves several rounds of conversations between human participants and the system, such a study requires many human participants and time to complete. For instance, the user study of Calendar.help (Cranshaw et al., 2017) lasted for 18 months. We believe that enabling a user study to be completed quickly would be beneficial in iterating over and expediting system designs.
Thus, to address these challenges, we alternatively conducted the evaluation study under synthetic meeting scheduling settings and used LLMs to simulate members’ conversation behavior. This approach enables more controlled and scalable experimentation while preserving a realistic level of human-like interaction.
@@ 5.1. Synthetic Data
[…]
We synthesized a collection of meetings, each of which includes: (1) an organizer, (2) members of the meetings, (3) the meeting subject, (4) date, and (5) duration. Example meeting data can be found in Table 5.
@@
[…]
Finally, the database $\mathcal{D}$ for this synthetic company encompasses all the public information about employees. The database includes the social graph and the profile data of all employees except for their schedule preferences (which are their private information and should be elicited by the system).
In the next section, we detail our experiment setup, evaluation metrics, and results.
@@ 5.2. Experiment Setup
We evaluated the performance of our system on the following axes: (i) the effect of the parameters of the system ($n,K$, incorporation of the database) on the performance, (ii) system performance compared to two baselines, (iii) robustness of our system subject to paraphrased preferences.
@@
  • Schedule preferences only. First, the system was required to propose meeting options based solely on the employees’ schedule preferences without considering the database $\mathcal{D}$. Namely, the system focused on finding a meeting time that balances individual schedule preferences elicited from meeting members. For this setting, we examined the effect of the meeting size $n$ and the number of proposed options $K$ as follows:
@@
  • Schedule preferences and database. Second, the coordination module considered not only schedule preferences but also the information in the database $\mathcal{D}$; the system also accounted for attendance constraints (if any) posed by the meeting organizer.
[…]
In all of the simulated meeting scenarios, the coordination processes were carried out over a total of $T=4$ rounds. This is consistent with the finding of Calendar.help (Cranshaw et al., 2017), which shows that most users had no more than five rounds of interactions with the system.
@@
In total, we ran 120 meeting scenarios. We note that throughout the two studies, we chose $n>K$, i.e., the meeting size was always greater than the number of proposed options. When there are more options presented at each round than the number of meeting members, it is difficult to amass a majority on one of the options and, therefore, reach a consensus. Moreover, the coordination process is designed with the goal of collecting the members’ feedback and gradually increasing the fraction of the members that the decision candidate can satisfy (i.e., satisfaction ratio $\hat{r}^{t}$). Therefore, the decision candidate is refined via several rounds of proposing new options, and proposing many options at each round is not needed.
Comparison with baselines. Our system includes two interactive elements: (i) the conversation between a user and the intent extraction module and (ii) the multiple rounds of discussion. To show the gains from the two elements, we compared the system with the following two baselines:
  • Single-Round Non-Conversational Baseline. In this baseline, both elements were absent. We did not use the interactive/conversational intent extraction module and instead provided preferences without processing them directly in the coordination module. We ran the protocol for only one round (i.e., $T=1$).
  • Single-Round Conversational Baseline. In this baseline, we used the interactive intent extraction module for the conversation with users but ran the system for only one round. This is to measure the extent to which multiple rounds of refining the decision candidate improves the system performance.
@@ 6. Simulation Study Results
In this section, we present the results of our study in Section 5. Figures 6, 4 and 5 summarize the evaluation metrics in the experiment of Section 5, which are plotted in the following manner: Given a meeting scenario, for each metric we calculated its percentage change at round $t$ with respect to its value at round $t=1$. Then we plotted the average of the percentage change in the metrics for each round, together with their 95% confidence intervals in our system [^9] We use the Student-t distribution with 19 degrees of freedom and a two-sided confidence interval.. The plots showing the metrics values for all scenarios are relegated to Appendix C.
[…]
@@ 6.1. Effect of Meeting Size $n$
[…]
In particular, the system performed significantly better in these three dimensions in the setting with $n=5$ members than in the ones with $n=3$ and $n=4$. For $n=5$ members, the average number of interactions was largely reduced by $\approx 50\%$ [^10] We use $\approx$” to for “approximately”., compared to $\approx 25\%$ ($n=3$) and $\approx 45\%$ ($n=4$). Secondly, the satisfaction score of the decision candidate for $n=5$ was eventually increased by $\approx 120\%$, while it was increased by $\approx 45\%$ ($n=3$) and $\approx 40\%$ ($n=4$) respectively. Similarly, the satisfaction ratio of the decision candidate was increased by $\approx 110\%$ for $n=5$, while both $n=3$ and $n=4$ experienced an increase of $\approx 40\%$. Finally, the equity score of the decision candidate decreased for $n=5$, but it exhibited subtle fluctuations for $n=3$ and $n=4$.
@@ 6.2. Effect of Number of Options $K$
Figure 5 shows the results of the experiment for the settings when the number of options $K$ varies. We found that increasing the number of options from $K=2$ to $K=4$ decreased the final percentage change of the satisfaction ratio. Specifically, for $K=2$ options, the system achieved an $\approx 100\%$ increase in the satisfaction ratio at the final round, whereas adopting $K=3$ and $K=4$ options yielded an increase of $\approx 50\%$. We also found that the increase in satisfaction score was $\approx 100\%$ for $K=2$, compared to $\approx 50\%$ ($K=3$) and $\approx 100\%$ ($K=4$). The equity score experienced a smaller percentage decrease as $K$ was increased. These observations indicated that the system’s performance was negatively impacted as the number of options $K$ was increased.
Figure 6. The average of the percentage change in each metric with the number of rounds for $T=4,n=5,K=2$ and whether the database $\mathcal{D}$ is considered or not. For each setting, 20 different meeting scenarios were tested.
6.3. Effect of Incorporating the database
Figure 6 compares the metrics between including and excluding the database for $T=4$, $n=5$ and $K=2$. When the database and attendance constraints were incorporated, the average number of interactions remained decreasing, and the satisfaction ratio and the decision candidate’s score increased. However, compared to using the schedule preferences only, the evaluation metrics deteriorated, and the system’s performance worsened in each dimension.